OpenStudy (anonymous):

A plane landing on a small tropical island has just 56 m of runway on which to stop. If its initial speed is 62 m/s, what is the maximum acceleration of the plane during landing, assuming it to be constant? Answer in m/s^2 How long does it take for the plane to stop with this acceleration? Answer in units of s

OpenStudy (anonymous):

what is the smallest acceleration required to stop

OpenStudy (anonymous):

Ok let's work on this one together

OpenStudy (anonymous):

We're assuming it is able to stop, for the smallest possible acceleration you will use the entire runway of 56 m, does that make sense?

OpenStudy (anonymous):

the formula i was using is v^2=v(initial)^2 +2a(x-x(initial)) and, yeah, that makes sense

OpenStudy (anonymous):

So ok, you have final velocity of 0, since you stopped, and initial velocity of 62, you also have distance travelled of 56m, I would recommend using this formula: Vf^2 = Vi^2 + 2ad Where Vf is final velocity, Vi is initial velocity, a is acceleration (you're solving for this) and d is the distance travelled

OpenStudy (anonymous):

So let's substitute your numbers into the formula: (0)^2 = (62) ^2 + 2*(a)*(56) From this I get a = -34.32 m/s^2

OpenStudy (anonymous):

awesome, i can see where i went wrong. thanks

OpenStudy (anonymous):

i can finish the next part

OpenStudy (anonymous):

There are really just 3 formulas for you to know for this topic: Vf^2 = Vi^2 + 2ad Vf = Vi + at d = Vi*t + 0.5*a*t^2

OpenStudy (anonymous):

k. thanks

OpenStudy (anonymous):

How long does it take for the plane to stop with this acceleration? Answer in units of s